# ROS2WASM: Bringing the Robot Operating System to the Web

### v1

@@ Abstract
The Robot Operating System (ROS) has become the de facto standard middleware in robotics, widely adopted across domains ranging from education to industrial applications. The RoboStack distribution has extended ROS’s accessibility by facilitating installation across all major operating systems and architectures, integrating seamlessly with scientific tools such as PyTorch and Open3D. This paper presents ROS2WASM, a novel integration of RoboStack with WebAssembly, enabling the execution of ROS 2 and its associated software directly within web browsers, without requiring local installations. This approach significantly enhances reproducibility and shareability of research, lowers barriers to robotics education, and leverages WebAssembly’s robust security framework to protect against malicious code. We detail our methodology for cross-compiling ROS 2 packages into WebAssembly, the development of a specialized middleware for ROS 2 communication within browsers, and the implementation of a web platform available at www.ros2wasm.dev that allows users to interact with ROS 2 environments. Additionally, we extend support to the Robotics Toolbox for Python and adapt its Swift simulator for browser compatibility. Our work paves the way for unprecedented accessibility in robotics, offering scalable, secure, and reproducible environments that have the potential to transform educational and research paradigms.
@@ I Introduction
Figure 1: Execution of a ROS 2 publisher and subscriber setup within a Jupyter Notebook environment running entirely in the browser, with no local installation required (https://mbatc.github.io/wasm-cpp-test/lab/). The notebook creates publishers for topics topic_a and topic_b, as well as two minimal subscribers listening to these topics. The code snippet at the bottom shows the concurrent execution of these nodes, with messages being published and received, verifying seamless communication between ROS 2 nodes directly within the browser-based interface.
@@
Replicating workspaces is particularly beneficial in scenarios such as teaching robotics courses [5] and reproducing research papers [6]. Instructors often struggle to provide students with the necessary tools, resorting to maintaining a computer lab or providing instructions for workplace replication. However, students’ diverse computing environments can lead to inconsistencies in how ROS packages function. ROS2WASM mitigates these issues by enabling students to run complex robotics simulations directly in their browsers, eliminating the need for local setup or reliance on university servers (Figure 1).
In research, reproducibility faces similar obstacles. Traditional environments are often rigid and difficult to adapt, wand version compatibility issues can impede reproducibility efforts. While Docker containers provide a potential solution, installing Docker and managing containers can be daunting and time-consuming for many aspiring roboticists [7, 8].
We propose a solution based on WebAssembly (WASM) [9, 10], a low-level bytecode format designed to run efficiently in web browsers. WASM is fast, secure, and portable, making it an ideal choice for running high-performance applications in the browser. It is well-sandboxed and supported by all major browsers, making it a universal target for web applications.
@@
ROS2WASM builds upon RoboStack [3], which integrates ROS with the Conda package manager. RoboStack offers several key advantages, including reproducible environments, robust version management, compatibility across platforms and Linux distributions, distribution of pre-built binaries, environment isolation, and simplified package creation.
@@
    A custom ROS 2 middleware implementation, rmw-wasm, cross-compiled into WebAssembly modules, enabling communication between ROS 2 modules within browser environments.
@@ II Related Works
In recent years, the robotics community has witnessed significant advancements in tools and frameworks that enhance the accessibility and functionality of the Robot Operating System (ROS). These efforts aim to make ROS more adaptable to various platforms and environments, including web-based applications. This section provides an overview of existing technologies related to WebAssembly, ROS in the browser, and browser-based robotics environments, highlighting how ROS2WASM advances the current state of the art.
@@ II-A Overview of ROS 2 and RoboStack
ROS 2 [1], the successor of ROS, builds on the strengths of its predecessor while addressing several limitations. It introduces a new communication middleware based on the Data Distribution Service (DDS), which provides real-time communication capabilities, improved security, and better support for multi-robot systems. DDS enables ROS 2 nodes to communicate over a network in a decentralized and scalable manner, making it ideal for distributed robotic applications. By leveraging DDS, ROS 2 ensures that data is exchanged reliably and efficiently between nodes, even in large, complex systems.
RoboStack [3] is a project that integrates ROS with the Conda package manager, significantly enhancing the accessibility and usability of ROS across different platforms. By leveraging Conda, RoboStack allows users to install ROS distributions in isolated environments, ensuring reproducibility and compatibility across different operating systems and architectures. This approach simplifies the installation process, reduces dependency conflicts, and allows for easier management of ROS packages. RoboStack also facilitates the distribution of pre-built binaries, which can be cross-compiled for various platforms. This capability is central to the development of our proposed ROS2WASM. By cross-compiling ROS to WebAssembly, we enable seamless deployment of ROS packages in a web environment.
@@ II-B WebAssembly
WebAssembly (WASM) is a relatively new technology in the web development landscape [9]. Released in 2017, it emerged from a collaboration between major browser vendors – Google Chrome, Microsoft Edge, Mozilla Firefox, and Apple Safari – following the completion of a Minimum Viable Product.
@@
Early demonstrations of WASM’s potential included porting the game AngryBots, developed by Unity Technologies, to the web [14]. Similarly, the Doom 3 engine was ported to WASM using Emscripten [15], showcasing the ability to run large, demanding C++ programs within a browser. WASM’s adaptability and performance make it an appealing choice for robotics applications.
@@ II-C ROS in the Browser
Another relevant project is the roswasm suite [19], an open-source collection of libraries designed to assist in cross-compiling C++ ROS nodes to WebAssembly. Despite its active maintenance, the roswasm suite is primarily focused on the outdated ROS 1 distribution, limiting its applicability to ROS 2 users.
@@ II-D Browser-Based Robotics Environments
Similarly, the rosbridge suite [22] offers packages that enable interaction with ROS systems via a JavaScript Object Notation (JSON) interface, allowing non-ROS systems to communicate with ROS by sending JSON messages. The suite supports both ROS 1 and ROS 2. Despite its extensive capabilities, rosbridge relies on an existing ROS installation and is therefore not a complete alternative for running ROS in the browser. The ROS Control Center [23] is a web application that uses rosbridge to establish a WebSocket connection to a robotics system running ROS. In contrast, ROSboard [24] does not depend on the rosbridge suite. Instead, it implements a custom Tornado web server that functions as both a web server and a WebSocket server, providing a lightweight and streamlined interface for interacting with ROS systems.
@@ III Methods and Demonstrations
This section details the methodology employed to develop ROS2WASM. We begin with the design of a custom middleware architecture that integrates ROS 2 with WebAssembly, enabling seamless communication within browser environments (Section III-A). Next, we describe the cross-compilation process, which extends the RoboStack ecosystem to support WebAssembly, ensuring that ROS packages can be built and deployed for web-based use (Section III-B). We then explain the web integration strategy that allows these packages to be served on a website using JupyterLite, facilitating easy access and execution within a browser (Section III-C). Additionally, we cover the adaptations made to the Robotics Toolbox for Python and the Swift Robotics Simulator to overcome the constraints of the sandboxed web environment, enabling these tools to function effectively in a browser (Section III-D). Finally, we discuss methods for establishing communication with non-ROS physical robots, demonstrating one exemplary practical application of ROS2WASM in real-world scenarios (Section III-E).
@@ III-A Custom Middleware Architecture
The middleware implementation is divided into two packages: wasm-cpp (implemented in C++) and wasm-js (implemented in JavaScript). Communication between these two packages is facilitated by emscripten::val, which allows JavaScript functions to be called from C++ and vice versa [16].
@@
Publishers and subscribers are the simplest participant types, responsible for sending and receiving messages, respectively. Service participants are derived from these basic types; service clients and servers consist of a publisher-subscriber pair (Figure 5). Before a service client can invoke a service, it verifies the availability of a service server through a query to wasm-js, which tracks all available entities. The client publishes a request to the service request topic and awaits the server’s response via the service response topic. The server publishes and subscribes to these topics in reverse. Figure 4 illustrates the relationships between the participants as implemented in the wasm-cpp package.
@@ III-B Cross-compilation
Several dependencies for ROS packages were not yet available for WebAssembly, necessitating the addition of patches tailored to the WASM platform. These patches often involved introducting preprocessor guards for Emscripten and are intended to guide future users on best practices and common pitfalls when targeting WebAssembly. Additionally, new Emscripten-specific compile flags were developed and applied during the package build process.
The overall build process closely follows that of RoboStack, with continuous integration set up to build and upload packages to an anaconda.org channel, available at https://github.com/mbatc/ros-humble.
@@ III-C Web Integration
After building individual packages, the next challenge is serving them on a website. This is facilitated by the JupyterLite build tool. By specifying the required packages in a deps.yaml file and running jupyter lite build, a complete website is generated, built on top of JupyterLite, with all specified packages installed. Users can then run commands such as import rclpy in a Python environment within the browser, using rclpy as they would in a typical ROS 2 setup (Figure 1). For those preferring C++, the rclcpp package can be used with the xeus-cpp kernel for JupyterLite.
@@ III-D Robotics Toolbox for Python and Swift
Adapting the Swift Robotics Simulator [27] to the sandboxed web environment required addressing limitations related to threading, networking, and file system access. We cross-compiled the simulator to WebAssembly using Emscripten Forge and the PyJS framework, making significant modifications to key components of Swift’s architecture. The Swift Simulator running in WASM demonstrates a method of adapting an existing project to work within the limitations of Emscripten and WASM, and the approach used here may inform future ports targeting Web Assembly. We successfully cross-compiled the simulator, and users can interact with it via a Python terminal running within a Web browser. The resulting version of Swift is hosted at https://mbatc.github.io/swift-wasm-test/, allowing users to run and interact with the simulator via a Python terminal embedded in a web page (Figure 7).
@@
Figure 7: Swift is a lightweight browser-based simulator built on top of the Robotics Toolbox for Python. This simulator provides robotics-specific functionality for rapid prototyping of algorithms, research, and education. Here, we introduce Swift running in the browser without any installation required: https://mbatc.github.io/swift-wasm-test/
Three major modifications were required to key components of Swift’s architecture. First, we needed a new messaging system to replace WebSockets. This was achieved by passing messages via Python/Javascript interoperability using PyJS. Second, we had to find an alternative way to launch the Swift environment without starting an HTTP server. We eliminated the HTTP server by loading the web interface in an embedded iframe. Third, we neeeded to add hooks to load Robotics Toolbox assets from the packaged Python environment. This was achieved by intercepting asset fetching so that assets could be retrieved from the PyJS file system.
@@ III-E Communication with a non-ROS Robot
To showcase the wide range of possible applications of ROS2WASM, we demonstrate the manipulation of a physical robot. Data can be sent and received from a robot with the publishers and subscribers developed in Section III-A, as long as there is a communication bridge between the robot and the web browser. There are two primary ways to communicate with a robot. The first way is to configure a robot to launch a rosbridge server to send information to the browser; however, this method requires the robot to be running ROS for creating a rosbridge.
An alternative way for robots which do not run ROS natively is to create an interface between the browser and the robot’s system. The bridge can be established via Bluetooth, WiFi, or a wired connection. Here, we demonstrate ROS2WASM with the LEGO BOOST Vernie robot (Figure 8). Vernie can be connected to any computer via Bluetooth. A pre-existing library, lego-boost-browser [28], to establish the interface between ROS and Vernie. This library uses the Web Bluetooth API to control the LEGO BOOST robot.
@@
The lego-boost-browser library can send commands to Vernie to change the LED color or activate the two motors for locomotion. We created a publisher that simply publishes different LED colors as strings from Vernie’s acceptable list of colors. When these messages reach the message stacks of the respective topic, they are forwarded to the robot through lego-boost-browser. Figure 8 shows how a web page running an LED color publisher appears. The user must connect to Vernie via Bluetooth before being able to send the commands to change LED color based on the last published message.
To control the position of the Vernie robot, we created a publisher that sends geometry_msgs/Twist messages. These messages contain linear and angular velocity vectors, allowing us to express the robot’s velocity in free space.
@@ IV Conclusion
We have demonstrated the feasibility and potential of running ROS 2 within a web browser using WebAssembly. Our work encompasses the design and implementation of three major components: a custom middleware architecture, a method for cross-compiling ROS 2 packages to WebAssembly, and the deployment of a web-based platform that allows users to interact with ROS environments directly in their browsers without the need for local installation or configuration.
Our website, www.ros2wasm.dev, enables the public to engage with a range of demonstrations that showcase the capabilities of ROS2WASM. These demonstrations vary in user interaction level and complexity, highlighting the versatility and scalability of the platform. By successfully running and facilitating communication between multiple ROS nodes within the browser, we have validated the core concept of ROS2WASM. This increased accessibility opens up new possibilities for education, research, and collaboration in robotics, making advanced tools more widely available. While the current platform allows easy sharing of these environments via a simple link, further efforts are needed to simplify the setup process for novice users, particularly in creating and deploying their own ROS-based web applications.
Looking ahead, future work could explore integrating Zenoh as the middleware for ROS 2 [29]. Zenoh, recently added as an official RMW implementation for ROS 2, offers the potential to natively support WebAssembly. This could eliminate the need for the custom middleware introduced in this project, streamlining the architecture and potentially improving performance and compatibility. Another promising direction is the adaptation of graphical user interfaces such as the popular rviz and MoveIt for browser compatibility. Such enhancements would further solidify the viability of running ROS 2 in the browser and could lead to even broader adoption and innovation in web-based robotics.

### v2

@@ Abstract
The Robot Operating System (ROS) has become the de facto standard middleware in robotics, widely adopted across domains ranging from education to industrial applications. The RoboStack distribution, a conda-based packaging system for ROS, has extended ROS’s accessibility by facilitating installation across all major operating systems and architectures, integrating seamlessly with scientific tools such as PyTorch and Open3D. This paper presents ROS2WASM, a novel integration of RoboStack with WebAssembly, enabling the execution of ROS 2 and its associated software directly within web browsers, without requiring local installations. ROS2WASM significantly enhances the reproducibility and shareability of research, lowers barriers to robotics education, and leverages WebAssembly’s robust security framework to protect against malicious code. We detail our methodology for cross-compiling ROS 2 packages into WebAssembly, the development of a specialized middleware for ROS 2 communication within browsers, and the implementation of www.ros2wasm.dev, a web platform enabling users to interact with ROS 2 environments. Additionally, we extend support to the Robotics Toolbox for Python and adapt its Swift simulator for browser compatibility. Our work paves the way for unprecedented accessibility in robotics, offering scalable, secure, and reproducible environments that have the potential to transform educational and research paradigms.
@@ I Introduction
Figure 1: Execution of a ROS 2 publisher and subscriber setup within a Jupyter Notebook environment running entirely in the browser, with no local installation required (www.ros2wasm.dev). The notebook creates publishers for topics topic_a and topic_b, as well as two minimal subscribers listening to these topics. The log output at the bottom shows the concurrent execution of these nodes, with messages being published and received, verifying seamless communication between ROS 2 nodes directly within the browser-based interface.
@@
Replicating workspaces is particularly beneficial in scenarios such as teaching robotics courses [5] and reproducing research papers [6]. Instructors often struggle to provide students with the necessary tools, resorting to maintaining a computer lab or providing instructions for workplace replication. However, students’ diverse computing environments can lead to inconsistencies in how ROS packages function due to differences in operating systems, library versions, and system configurations. ROS2WASM mitigates these issues by providing a standardized execution environment within the browser itself. By compiling ROS 2 to WebAssembly, students can interact with identical software configurations regardless of their hardware or operating system, running ROS 2 directly in their browsers (e.g., Chrome, Firefox, Safari, Edge). ROS2WASM thus eliminates any local setup or reliance on university servers (Figure 1).
In research, reproducibility faces similar obstacles. Traditional environments are often rigid and difficult to adapt, and version compatibility issues can impede reproducibility efforts. While Docker containers provide a potential solution, installing Docker and managing containers can be daunting and time-consuming for many aspiring roboticists [7, 8].
We propose a solution based on WebAssembly (WASM) [9, 10], a low-level bytecode format designed to run efficiently in web browsers. WASM is fast, secure, portable, and well-sandboxed across all major browsers, making it ideal for high-performance web applications.
@@
Although ROS2WASM involves initial complexity in cross-compilation and middleware development, this complexity is largely encapsulated on the developer side. For end users, ROS2WASM offers a lower barrier to entry compared to Docker and similar solutions. ROS2WASM builds upon RoboStack [3], which integrates ROS with the Conda package manager. RoboStack offers several key advantages, including reproducible environments, robust version management, compatibility across platforms and Linux distributions, distribution of pre-built binaries, environment isolation, and simplified package creation.
@@
    A custom ROS 2 middleware implementation, rmw-wasm, cross-compiled into WebAssembly modules, enabling communication between ROS 2 modules within browser environments. This middleware also enables communication with external physical robots running standard ROS 2 installations using https://wiki.ros.org/roslibjs.
@@ II Related Works
This section overviews existing technologies related to WebAssembly, ROS in the browser, and browser-based robotics environments that contextualize our work.
@@ II-A Overview of ROS 2 and RoboStack
ROS 2 [1], the successor of ROS, builds on the strengths of its predecessor while addressing several limitations. It introduces a new communication middleware based on the Data Distribution Service (DDS), which provides real-time communication capabilities, improved security, and better support for multi-robot systems. DDS enables ROS 2 nodes to communicate over a network in a decentralized and scalable manner, making it ideal for distributed robotic applications.
RoboStack [3] is a project that integrates ROS with the Conda package manager, significantly enhancing the accessibility and usability of ROS across different platforms. By leveraging Conda, RoboStack allows users to install ROS distributions in isolated environments, ensuring reproducibility and compatibility across different operating systems and architectures. This approach simplifies the installation process, reduces dependency conflicts, and allows for easier management of ROS packages. RoboStack also facilitates the distribution of pre-built binaries, which can be cross-compiled for various platforms. This capability is central to the development of our proposed ROS2WASM, where we cross-compile ROS to WebAssembly to enable seamless deployment of ROS packages in a web environment.
@@ II-B WebAssembly
WebAssembly (WASM) is a relatively new technology in the web development landscape [9], which emerged in 2017 based on a collaboration between major browser vendors – Chrome, Edge, Firefox, and Safari.
@@
Early demonstrations of WASM’s potential included porting the game AngryBots, developed by Unity Technologies, to the web [14]. Similarly, Doom 3 was ported to WASM using Emscripten [15], showcasing the ability to run large, demanding C++ programs within a browser. WASM’s adaptability and performance make it appealing for robotics applications.
@@ II-C ROS in the Browser
Another relevant project is the roswasm suite [19], an open-source collection of libraries designed to assist in cross-compiling C++ ROS nodes to WebAssembly. However, the roswasm suite is primarily focused on the outdated ROS 1 distribution, limiting its applicability to ROS 2 users.
@@ II-D Browser-Based Robotics Environments
The rosbridge suite [22] enables non-ROS systems to communicate with ROS via JSON, supporting both ROS 1 and ROS 2 but requiring existing ROS installations, thus not being a complete alternative for running ROS in the browser. The ROS Control Center [23] is a web application that uses rosbridge to establish a WebSocket connection to a robotics system running ROS. In contrast, ROSboard [24] does not depend on the rosbridge suite. Instead, it implements a custom Tornado web server that functions as both a web server and a WebSocket server, providing a lightweight and streamlined interface for interacting with ROS systems.
@@ III Methods and Demonstrations
This section details the methodology employed to develop ROS2WASM. We begin with the design of a custom middleware architecture for ROS 2 with WebAssembly, enabling communication within browser environments (Section III-A). Next, we describe the cross-compilation process extending RoboStack to support WebAssembly (Section III-B), followed by the web integration strategy using JupyterLite for browser-based access (Section III-C). Additionally, we cover adaptations to the Robotics Toolbox for Python and Swift Robotics Simulator for the sandboxed web environment (Section III-D). Finally, we discuss methods for communicating with non-ROS physical robots, demonstrating a practical application of ROS2WASM (Section III-E).
@@ III-A Custom Middleware Architecture
The middleware implementation is divided into wasm-cpp (implemented in C++) and wasm-js (implemented in JavaScript). Communication between these two packages is facilitated by emscripten::val, which allows JavaScript functions to be called from C++ and vice versa [16].
@@
Publishers and subscribers are the simplest participant types and are responsible for sending and receiving messages, respectively. Service participants are derived from these basic types; service clients and servers consist of a publisher-subscriber pair (Figure 5). Before a service client can invoke a service, it verifies the availability of a service server through a query to wasm-js, which tracks all available entities. The client publishes a request to the service request topic and awaits the server’s response via the service response topic. The server publishes and subscribes to these topics in reverse. Figure 4 illustrates the relationships between the participants as implemented in the wasm-cpp package.
@@ III-B Cross-compilation
All ROS 2 packages must be recompiled specifically for WebAssembly to function in the ROS2WASM environment. For external libraries and dependencies, our approach builds upon the https://emscripten-forge.org, which is planned to be incorporated into the conda-forge ecosystem in the future. Emscripten-forge already provides many commonly used external libraries (including Python packages and numerical toolboxes) pre-compiled for WebAssembly.
As part of the development efforts for ROS2WASM, several dependencies were added to emscripten-forge, which sometimes necessitated the addition of patches tailored to WASM, e.g. introducing preprocessor guards for Emscripten. These patches are intended to guide future users on best practices and common pitfalls when targeting WebAssembly.
The overall build process closely follows that of RoboStack, with continuous integration set up to build and upload packages to an anaconda.org channel, available at https://github.com/RoboStack/ros-humble.
@@ III-C Web Integration
After building individual packages, the next challenge is serving them on a website. Our approach enables two distinct types of web deployments: For applications compiled from C/C++ source code (as shown in Figure 2), we directly cross-compile the code to a WASM module and serve it as a standalone web application (Figure 8). For writing and executing ROS 2 applications directly in the browser, we utilize JupyterLite. By specifying the required packages in an environment.yaml file and running jupyter lite build, a complete website is generated, built on top of JupyterLite, with all specified packages installed. Users can then run commands such as import rclpy in a Python environment within the browser, using rclpy as they would in a typical ROS 2 setup (Figure 1). For those preferring C++, the rclcpp package can be used with the xeus-cpp kernel for JupyterLite.
@@ III-D Robotics Toolbox for Python and Swift
This section briefly reviews how non-ROS robotics applications can be adapted for WASM. While the Robotics Toolbox for Python required no changes, adapting the graphics-based Swift Robotics Simulator [27] to the sandboxed web environment meant addressing limitations in threading, networking, and file system access. We cross-compiled Swift to WebAssembly using Emscripten and the PyJS framework, which required three major modifications.
@@
Figure 7: Swift is a lightweight browser-based simulator built on top of the Robotics Toolbox for Python. This simulator provides robotics-specific functionality for rapid prototyping of algorithms, research, and education. Here, we introduce Swift running in the browser without any installation required: www.ros2wasm.dev.
First, we needed a new messaging system to replace WebSockets, which are not fully supported in WASM. This was achieved by passing messages via Python/Javascript interoperability using PyJS. Second, we had to find an alternative way to launch the Swift environment without starting an HTTP server. We eliminated the HTTP server by loading the web interface in an embedded iframe. Third, we needed to add hooks to load Robotics Toolbox assets from the packaged Python environment. This was achieved by intercepting asset fetching so that assets could be retrieved from the PyJS file system.
The resulting implementation demonstrates an approach for adapting existing projects to work within Emscripten and WASM constraints, which may inform future WebAssembly ports. Users can now interact with the simulator through a Python terminal embedded in a web page (Figure 7).
@@ III-E Communication with a non-ROS Robot
To showcase the wide range of possible applications of ROS2WASM, we demonstrate the manipulation of a physical robot. Data can be sent and received from a robot with the publishers and subscribers developed in Section III-A, as long as there is a communication bridge between the robot and the web browser. There are two primary ways to communicate with a robot: Firstly, configure a robot to launch a rosbridge server to send information to the browser; however, this method requires the robot to be running ROS for creating a rosbridge. An alternative way for robots which do not run ROS natively is to create a bridge between the browser and the robot’s system. The bridge can be established via Bluetooth, WiFi, or through a wired connection. Here, we demonstrate ROS2WASM with the LEGO BOOST Vernie robot that is connected to any computer via Bluetooth. A pre-existing library, lego-boost-browser [28], is used to establish the interface between ROS and Vernie via a Web Bluetooth API.
@@
The lego-boost-browser library can send commands to Vernie to change the LED color or activate the two motors for locomotion. We created a publisher that simply publishes different LED colors as strings from Vernie’s acceptable list of colors. When these messages reach the message stacks of the respective topic, they are forwarded to the robot through lego-boost-browser. Figure 8 shows how a web page running an LED color publisher appears, displaying the user interface with connect/disconnect buttons, publisher controls, and a console that logs the sequence of colors being published in real-time to demonstrate the robot’s response to commands. To control the position of the Vernie robot, we created a publisher that sends geometry_msgs/Twist messages that contain linear and angular velocity vectors to express the robot’s velocity in free space.
@@ IV Conclusion
We demonstrated the feasibility and potential of running ROS 2 within a web browser using WebAssembly. Our work encompasses the design and implementation of three major components: a custom middleware architecture, a method for cross-compiling ROS 2 packages to WebAssembly, and the deployment of a web-based platform that allows users to interact with ROS environments directly in their browsers without the need for local installation or configuration.
Our website, www.ros2wasm.dev, enables the public to engage with a range of demonstrations that showcase the capabilities of ROS2WASM. These demonstrations vary in user interaction level and complexity, highlighting the versatility and scalability of the platform. By successfully running and facilitating communication between multiple ROS nodes within the browser, we have validated the core concept of ROS2WASM. This increased accessibility opens up new possibilities for education, research, and collaboration in robotics, making advanced tools more widely available. The current platform allows easy sharing of these environments via a simple link. However, further efforts are needed to simplify the setup process for novice users, particularly in creating and deploying their own ROS-based web applications.
ROS2WASM also has significant potential in industrial applications. Web-based ROS could enable robust server-based distributed control systems for multi-agent coordination, allowing operators to monitor and control robot fleets from any device without specialized software installations. Additionally, ROS2WASM could facilitate scalable and flexible cloud-based robotic processing, where computationally intensive tasks can be executed on remote servers while maintaining responsive browser-based interfaces on local devices.
Looking ahead, future work could explore integrating Zenoh as the middleware for ROS 2 [29]. Zenoh, a lightweight communication protocol recently added as an official RMW implementation for ROS 2, offers the potential to natively support WebAssembly. This could eliminate the need for our custom middleware, thereby streamlining the architecture and further improving performance and compatibility.
Another promising direction is adapting packages that contain graphical user interfaces such as the popular rviz, Gazebo, and MoveIt for browser compatibility. Such adaptations would further solidify the viability of running ROS 2 in the browser, which could lead to even broader adoption and innovation in web-based robotics.